OpenStudy (anonymous):

help please Simplify (x – 3)(x2 + 7x – 8)

OpenStudy (muhammad_nauman_umair):

=x(x^2+7x-8)-3(x^2+7x-8) =x^3+7x^2-8x-3x^2-21x+24 =x^3+4x^2-29x+24

OpenStudy (anonymous):

thank you!